1.      A company has the following demand equation

 

Q = 1000 – 3000 P + 10 A

 

Q = Quantity demanded

 

P = Product Price

 

A = Advertisement expenditure

 

Assume that P = 3 and A = 2000

 

Ø Suppose the firm drops the price to Rs. 2.50 would this be beneficial.

Ø Suppose the firm raises the price to Rs. 4.00 while increasing its advertisement expenditure by 100 would this be beneficial? Explain



1
Expert's answer
2021-10-25T17:43:52-0400

a.

The demand equation is given as:

Q = 1000 - 3000P + 10A

Let P = 3 & A = 2000

If the firm drops the price to Rs. 2.50 then quantity demanded would be:

Qd1 = 1000 - 3000(3) + 10(2000) = 12,000

Qd2 = 1000 - 3000(2.5) + 10(2000) = 13,500

If the firm drops the price to Rs. 2.50 then it would be beneficial. It is because decrease in price would increase quantity demanded.


b.

If the firm raises the price to Rs. 4.00 while increasing its advertisement expenditure by 100 then quantity demanded would be:

Qd = 1000 - 3000(4) + 10(2100) = 10,000

The increase in price would not be beneficial for the firm due to adverse effect on quantity demanded.